# Evaluating Sound Similarity Metrics for Differentiable, Iterative Sound-Matching

### v1

@@ Abstract
Index Terms:
Audio synthesis, differentiable digital signal processing, music information retrieval, sound-matching
@@ I Introduction
I-A Digital Synthesis With Sound-Matching
Since the 1970s, artists such as Suzanne Ciani [1] have explored the use of audio synthesizers to replicate natural sounds—such as a bottle cap popping off—in ways that preserve the essential qualities of the original while showcasing the unique characteristics of the synthesizer. Rather than merely replicating sounds, this approach uses synthesizers for artistic reinterpretation. However, the automation of this interpretive approach to sound design using digital audio synthesizers has rarely been explored, and significant knowledge gaps remain even in the simpler problem of replicating sounds. One such gap is the lack of diversity and depth in the analysis of sound-similarity measures and their interaction with various synthesis methods. As a stepping stone toward automated interpretive sound design, we use automatic and manual sound-evaluations to investigate whether the optimal choice of a sound-similarity measure is dependent on the synthesis method.
A digital audio synthesizer is any software used for the creation and manipulation of audio. Since the advent of digital signal processing (DSP) in the 1960s [2], a wide variety of parametric audio generation functions have been proposed [3, 4, 5]. Manual sound design with a synthesizer is an iterative process of trial and error; when the artist has a desired (or target) sound in mind, sound design involves listening to the output, determining whether the output is closer or farther from the target, and adjusting the parameters of the synthesizer accordingly until the target is reached [4]. This paper explores the automation of this process, which we call iterative sound-matching.
Automatic sound design has been a subject of research for decades, with reduction of tinkering time and creation of “interesting” sounds as the main motivators [6, 7, 8, 9, 10, 11, 12, 5, 13, 14]. Sound-matching is a common form of automatic sound design explored in literature, where a similarity metric (or loss function) is given, and the goal is to find parameters for a synthesizer that replicate all or some of the characteristics of a target sound as best as possible [8, 12, 14]. Our analysis of past works shows that this general goal of “sound-matching” can be broken down in to very different subtypes, many of them rarely or never explored, despite their relevance to practical applications. This leads us to a typology of sound-matching tasks, and the importance of differentiable, iterative sound-matching, rarely explored in previous literature [15].
Iterative sound-matching refers to works that mimic the back-and-forth process of listening to a synthesizer’s output and updating the parameters until the termination conditions are met. The two necessary components are the loss function, which determines how close two sounds are, and the parametric synthesizer that outputs sounds based on input parameters. In the iterative workflow, a target sound is selected, the synthesizer outputs sounds based on its parameters, and the loss function determines the proximity of the output to the target. From there, the synthesizer parameters are updated, the output and target are compared, and so on.
Despite the early popularity of the iterative approach [16, 8, 12], recent works in sound-matching have favored supervised learning using inference models, deep neural-networks (DNNs), and large datasets. Rather than continually interacting with the environment, these models learn from a dataset of inputs (synthesizer parameters) and outputs (corresponding sound), and predict the synthesizer parameters for new sounds [17, 18, 19, 13, 14].
Iterative sound-matching is interesting for multiple reasons. First, it better resembles the human approach to sound-design, which requires continually interacting with the environment and making decisions according to a feedback signal [20]. Second, the trial-and-error approach leaves a trail of parameter sets from the starting point to the end goal, which could be of use to sound designers. Third, relative to the DNN-based inference models, it provides a more tractable environment with room for creative approaches to sound-similarity definitions, and more amenable to deployment on different synthesizers as it does not require pre-training on large datasets.
This work takes a differentiable approach to iterative sound-matching, a natural extension of early works in iterative sound-matching combined with recent advances in Machine Learning (ML) software [21, 22] and increase in computational power [23, 24]. A differentiable approach requires differentiable loss and synthesizer functions, allowing the synthesizer parameters to be updated using gradient descent [25, 26].
We discuss the history of sound-matching, and provide a nomenclature for different subtypes of sound-matching tasks. A number of weaknesses in current literature (see Section III-C) are identified. The experiments in this work focus on two of these core issues, referred to as Loss Selection and Synthesis Selection. The former refers to the search for an optimal loss function, lack of novel algorithms, and the—perhaps needless—focus on outperforming state of the art (SOTA), while the latter refers to the lack of diversity in synthesis methods.
In the case of Loss Selection, there have been many works and experiments comparing the accuracy of loss functions and choosing the best performers [15, 7, 11, 27, 18, 13, 7, 28]. We note that these claims regarding the effectiveness of one function versus another have consistently been made in limited contexts that may not generalize to other settings. One such limitation is the synthesizers used, and ties into the problem of Synthesis Selection, as typically only a single (and often, simple) synthesizer has been used for testing the effectiveness of loss functions. Furthermore, there is a lack of diversity in the underlying functions used by the synthesizers, with simple Frequency Modulation (FM) being the most common method of synthesis [8, 15, 12, 29].
Considering the interplay between the two problems, we design experiments that address them simultaneously. Testing the hypothesis that the performance of a similarity measure (or loss function) is influenced by other factors in the environment, particularly the method of synthesis. We combine 4 loss functions with 4 differentiable synthesizers, and compare their performance as measured by the similarity of the final output to the target. We evaluate the final similarity with two different automatic methods, as well as manual listening tests conducted by two of the authors. We compare the distributions of the evaluation scores in order to rank the loss functions from best to worst. The results of these experiments show that when accounting for different synthesizers, there is no consistent choice of sound similarity measure for optimal performance in sound-matching.
We note that for some synthesizers, our novel use of “dynamic time warping of envelopes” as a loss function for differentiable sound-matching outperforms the standard methods of audio comparison. These results imply that rather than aiming to outperform the SOTA and finding a universal best performer, more effort could be placed towards diversifying the existing similarity measures of sound as well as conducting experiments using a variety of synthesis methods. Sound design is an inherently creative endeavor, and would likely benefit from a diversified range of solutions for different approaches.
II Background
In this Section, we give a formal definition of sound-matching and the necessary components. These discussions lead to a nomenclature for types of sound-matching, as well as major perennial issues in the field.
II-A Digital Audio Synthesis
A digital audio synthesizer is any software used for the creation and manipulation of digital audio. Digital synthesizers use signal processing chains with a variety of digital signal processing (DSP) functions to create sounds. These functions are often parametric, and the set of parameters given to the synthesizer’s chain of DSP functions is called a synthesizer program.
The simplest form of DSP function is a sinusoidal tone. For example, consider:
$x(n)=sin(2\pi nT)$
where $T$ represents time in seconds, and $n$ is the frequency parameter. If $n$ is 1, then $x(n)$ would be a waveform with a frequency of 1 hertz, meaning that the value of $x(n)$ oscillates to the same point once every second. Waveform generators are called oscillators for this reason.
@@ II-A Digital Audio Synthesis
Since the advent of digital signal processing in the 1960s [2], a wide variety of parametric audio generation functions have been proposed [3, 4, 5]. This includes oscillators, filters, equalizers, and envelopes, which can be used sequentially or in parallel [3, 4]. Sound design with a synthesizer is done by modifying the parameters of the functions until reaching a desired output [30, 31].
A relatively small subset of these synthesis methods have been used for automatic sound-matching, notably additive and/or subtractive synthesizers [10, 17, 12, 9], physical modeling [32], and frequency/amplitude modulation (FM/AM) [8, 15].
Differential DSP (DDSP) is a general term for applications that combine machine learning techniques such as gradient descent [25, 26] with DSP [11]. Implementing complex DSP functions in a differentiable manner can be challenging, and effective differentiable audio similarity measures require careful mathematical expression of the desired attributes of sound. These issues have likely contributed to the limited exploration of this domain. Perhaps to most relevant recent work is by Vahidi et al., featuring the comparison of two differentiable loss functions with a simple FM synthesizer [15].
II-B Sound Representation and Loss Functions
@@
$L(\theta,t)=d\langle\phi(t),\phi(x)\rangle$
A proxy representation is the output of the function $\phi$, which can be thought of as a feature extraction function that maps the sounds $t$ and $x$ to their respective representations. The proportionality or distance metric $d$ has typically been the L1 distance [7, 35], calculated as the mean of the absolute difference between every point in the proxy representation [11, 15]:
$L(\theta,t)=\left\|\phi(t)-\phi(x)\right\|_{1}$
Here we discuss four methods of audio representation and the corresponding loss functions. Three of which have been used in previous works, and a novel method not previously used in iterative sound-matching.
II-B1 Parameter Loss
@@ II-B2 Fourier Spectrograms
Fourier-based transformations such as short-time Fourier transforms (STFT), Mel-spectrograms, and Mel-frequency cepstral coefficients have been viewed as the de facto and state-of-the-art representation of audio [36, 12, 17], however, there are many issues associated with their use in sound-matching [7, 15, 18, 27]. Fourier transformations allow for the conversion of a signal from the time-domain to the frequency domain. Audio spectrograms can be generated by segmentation of a piece of audio into overlapping windows followed by the application of Fourier transforms to each window. They are costly to compute, but provide a better temporal view of changes in frequency content [37, 34]. There are different types of spectrograms that have a basis in Fourier transformations, but the most notable and commonly used is the STFT.
What we call Fourier-based Spectrograms are variations on the STFT approach. For example, Mel-Spectrograms bin frequencies on a near-logarithmic scale to better match human perception of frequencies [37]. Multi-scale spectrograms (MSS) used in recent works are a simple weighted average of multiple spectrograms with different parameters such as window size, number of frequency bins, and hop size [11, 15]; this may provide some improvements at a higher computational cost [7, 11].
II-B3 Joint-Time Frequency Spectrum
@@
Recent works have focused on the limitations of parameter and spectral loss functions in sound-matching [15, 27], seeking to create more effective general solutions for the comparison of audio. Noting the aforementioned weaknesses of comparing STFT spectrograms, Vahidi et al. proposed differentiable Joint-Time Frequency Scattering (JTFS) [38] as an alternative to spectrogram loss in sound-matching, and showed improved performance in sound-matching with differentiable chirplet synthesizers [15]. JTFS is the result of the application of a 2D wavelet transformation to the time-frequency representation of a signal [38].
II-B4 Dynamic Envelope Warping
@@
Dynamic Time Warping (DTW) is a method for measuring similarity between multi-dimensional time-series [39, 37, 40]. Given any two time-series $X=\{x_{1},x_{2},...,x_{m}\}$ and $Y=\{y_{1},y_{2},...,y_{n}\}$, we have indices $i\in\{1...m\}$ and $j\in\{1...n\}$ defining $X$ and $Y$. When the series are warped, these indices change to expand or contract different portions of the series. To borrow the notation given by Muller [37], warped indices are a sequence $p=(p_{1},...,p_{L})$, where $p_{\ell}=(m_{\ell},n_{\ell})\in[1:m]\times[1:n]\text{ for }\ell\in[1:L]$, meaning that the indices for $X$ and $Y$ are reorganized under special conditions. In classical DTW, these conditions are monotonicity, boundary matching, and single step-size.
DTW measures the distance between the time-series after alignments, typically using Euclidean distance. The distance between a time-series and shifted versions of itself would be 0, regardless of shift amount [41]. Additional rules can be imposed to keep alignments locally constrained [42, 43].
II-C Automatic Sound Matching
In past works the sound-matching problem has been tackled from a variety of perspectives as it involves a number of modular parts: the choice of synthesizer $g$, the target sounds of interest, the representation function $\phi$, and the heuristic for finding the optimal $\theta$. To give a formal definition of sound-matching, we expand the definition given by recent works [15, 18]. The major components are as follows:
  • $g(\theta)$: Parametric audio synthesizer $g$ which takes on parameters $\theta$
  • $x$: The output of $g$, given a set of parameters $\theta$ or $g(\theta)=x$
  • $t$: The target sound which we want to replicate or imitate.
  • $\phi$: Representation function or feature extractor. $\phi$ is applied to $x$ and $t$ to facilitate their comparison by the loss function.
  • $L$: Loss or error function. $L(\theta,t)$ is a measure of distance between $x$ and $t$. Often proportional to the subtraction of their representations, or $\phi(x)-\phi(t)$
It could be argued that the implementation of these components is not an optimization problem with a correct answer, but a creative endeavor depending on the artist’s needs. Consider the hypothetical case where the goal might be to create an 8-bit [44] version of a high quality, organic snare drum: in such a case, there is no optimal or “correct” version of an 8-bit snare sound.
II-D Sound Matching Types
Sound-matching tasks can be categorized based on the choices of target sound $t$, the loss $L$ and representation function $\phi$, and the heuristic used for minimizing $\theta$. We will discuss the subtypes of sound-matching tasks in the upcoming sub-sections.
II-D1 In-Domain vs Out-of-Domain
@@
The choice of domain depends on whether we want to use the same synthesizer for the target and output sounds, which we call in-domain, or have target sounds that came from sources other than the synthesizer, or out-of-domain. To paraphrase the description given by Masuda et al. [13], if $g$, the synthesizer of choice, can accurately replicate the target sound $t$, or put differently, if $t$ itself is an output of $g$, then the sound-matching task is in-domain. If $t$ is not an output of $g$, then the sound-matching task is out-of-domain. In-domain tasks in general are simpler, and often there is a guarantee that there is a correct answer to the sound-matching problem, particularly if the goal is accurate replication of the sound. If the target sound is out-of-domain, replication is not guaranteed, and the goal becomes the imitation of some aspect of sound.
II-D2 Replication vs Imitation
@@
II-D3 Supervised Versus Direct Optimization
The goal of sound-matching is to find the optimal parameters $\theta^{*}$ that minimize the loss between the synthesizer output and the target sound.
$\theta^{*}=\arg\min_{\theta}L(\theta,t)$
The heuristics in past works can be broadly split into two categories:
  1. 1.
    Supervised (or inference) methods which use datasets of sound and parameter pairs to describe the generation objective. Often with a large number of input and output examples [11, 9, 17, 10]
  2. 2.
    Unsupervised or direct sound optimization methods [8, 12, 17, 15]. Usually require at most one target example, and a loss function that measures the difference between the target and synthesizer outputs.
Genetic algorithms (GA) [45] can be used for direct sound optimization, and have been the earliest and most common heuristic in sound-matching [8, 12, 17]. These algorithms start with arbitrary parameter sets that can be treated as an evolving population where the genomes are the parameter values. The most fit members of the group are the parameters which perform the best in the loss function, and create a new generation of parameters via mutation (random change in subset of parameters) and crossovers (combination of parameter sets); this process repeats until the goal or a maximum number of generations is reached.
Differentiable settings have the benefit of making goal-oriented updates to the synthesizer parameters (the goal being the minimization of loss). There are also drawbacks: other than requiring more computation power, differentiable functions require careful implementation of operations that are continuous and numerically stable; this makes the implementation of signal processing functions quite difficult, contributing to the scarcity of works in this domain.
II-E Iterative Sound Matching
$g(\hat{\theta})=x$$L(\hat{\theta},t){\ \propto\ }\phi(x)-\phi(t)$Optimize $\hat{\theta}$ew $\hat{\theta}$tartRandom $\hat{\theta}$
Figure 1: Iterative approach to sound design.
In the upcoming “Past Works” section, we will discuss the evolution of prior sound-matching research, with differentiable iterative sound-matching being the latest extension of the field.

### v2

@@ I Introduction
A digital audio synthesizer is any software used for the creation and manipulation of audio. The possibilities afforded by these synthesizers are infinite, leading to their widespread adoption by artists and sound-designers [1, 2, 3]. A typical synthesizer has a number of parameterizable functions which affect the sound output in various ways, and manual sound-design often involves modifying the parameters until a conceptualized target sound is reached. The automation of this approach has commonly been referred to as ”sound-matching”, with reduction of tinkering time and creation of “interesting” sounds as the main motivators  [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. The main requirements of sound-matching are a target sound, a similarity metric (or loss function), and a heuristic to find parameters for a synthesizer that replicate all or some of the characteristics of the target sound as best as possible [6, 10, 13]. However, despite the seeming simplicity of the problem and decades of research, sound-matching is not yet a practical solution for sound-designers. Here we are motivated to understand why this is the case, and what future research should prioritize in order to improve the field.
We provide a review of past works and identify major perennial issues in the field. In particular, we are concerned with the issues of Loss Selection and Synthesis Selection. The former refers to the search for an optimal loss function, lack of novel algorithms, and the—perhaps needless—focus on outperforming state of the art (SOTA), while the latter refers to the lack of diversity in synthesis methods. As a consequence of these problems, we find that not much is known about the interaction between different methods of synthesis and different loss functions. Is there a universally best performing loss function for audio? Is there a need for further development of bespoke loss functions?
The main hypothesis here is that the performance of a similarity measure (or loss function) is influenced by other factors in the environment, particularly the method of synthesis. Testing this hypothesis requires a variety of sound-matching experiments that measure whether a single loss function proves most effective. While there have been many works and experiments comparing the accuracy of loss functions [14, 5, 9, 15, 16, 12, 5, 17], we note that claims regarding the effectiveness of one function versus another have often been made in limited contexts that may not generalize to other settings. Beyond this central hypothesis, we also investigate three additional research questions: (Q1) To what extent do automatic evaluation metrics agree with manual listening tests? (Q2) Can loss functions based on Dynamic Time Warping (DTW) and Scale-Invariant Mean Squared Error (SIMSE) provide advantages over SOTA loss functions, and under what conditions? (Q3) Is iterative differentiable optimization a viable strategy for design of sound-matching experiments?
We adopt a lesser used iterative and differentiable approach to defining sound-matching experiments. The iterative approach better mimics the manual process of recursive listening and parameter adjustments towards sound design, while a differentiable environment allows access to the loss function gradients that can be used to better understand the nature of the problem. We define four differentiable synthesizers (each showcasing a fundamental method of synthesis used in modern synthesizers) and pair them with four different loss functions (two established methods, one utilizing DTW, and one utilizing SIMSE). We evaluate the final similarity with two different automatic methods, as well as manual listening tests conducted by two of the authors. We apply statistical ranking to the evaluation scores to compare outcomes across synthesizer–loss combinations and measure the similarity of different evaluation measures.
Contributions. The contributions of this paper include (1) an evaluation of multiple differentiable losses across multiple synthesis methods, (2) the introduction and justification of loss functions utilizing DTW and SIMSE (3) a discussion on the utility and agreement between manual and automatic evaluation metrics (4) a nomenclature of different sound-matching approaches and unsolved issues in the field.
II Background and Related Work
Sound-matching sits at the intersection of digital signal processing, audio representation, and optimization. Here we first formalize the sound-matching task, then review core synthesis methods, similarity measures, and optimization approaches. We conclude with a survey of related work and a discussion of gaps in the field.
II-A Formalization of Sound-Matching
Following prior works [14, 16], we define sound-matching in terms of a parametric synthesizer, a target sound, a representation function, and a similarity measure. The key components are:
  • $g(\theta)$: A parametric audio synthesizer $g$ with parameters $\theta$.
  • $x_{\theta}$: The synthesizer output, $x_{\theta}=g(\theta)$.”.
  • $x_{0}$: The target sound to be replicated or imitated.
  • $\phi(\cdot)$: A representation (feature extraction) function mapping signals into a comparison space.
  • $L$: A loss function that measures distance between $x_{\theta}$ and $x_{0}$, typically via $L(\theta,x_{0})=d(\phi(x_{\theta}),\phi(x_{0}))$ for some metric $d$.
This formalization highlights that sound-matching is not a single well-defined optimization problem, but rather depends on design choices for $g$, $\phi$, and $L$. Moreover, the “correct” solution may not exist in a strict sense: depending on the artistic goal, multiple parameter settings may yield acceptable or even preferable outputs. In the following sections we will discuss the various components of sound-matching, the subjectivity of “correct” solutions, and important issues in the field.
$g(\hat{\theta})=x_{\theta}$$L(\hat{\theta},x_{0})\propto\phi(x_{\theta})-\phi(x_{0})$Optimize $\hat{\theta}$New $\hat{\theta}$StartRandom $\hat{\theta}$
Figure 1: Iterative approach to sound design.
II-B Digital Signal Processing and Synthesis
A digital audio synthesizer generates or processes audio by chaining digital signal processing (DSP) functions. Each function is parameterized, and the set of parameters defining a chain of DSP functions is called a synthesizer program.
The simplest DSP function is a sinusoidal oscillator:
$x[n]=\sin\!\left(2\pi f\frac{n}{SR}\right),$
where $f$ is frequency in hertz, $SR$ is the sampling rate, and $n$ is the discrete time index. At a sampling rate of $SR$ samples per second, a 1 Hz oscillator completes one cycle per second. Such computations form the foundation of digital audio synthesis [1].
@@ II-B Digital Signal Processing and Synthesis
Since the advent of DSP in the 1960s [3], a wide variety of parametric functions have been proposed, including oscillators, filters, equalizers, and envelopes [1, 2]. Sound design involves modifying these parameters until a desired output is reached [18, 19].
Early sound-matching research focused heavily on frequency and amplitude modulation (FM/AM) synthesis [6, 10, 14], which is simple to implement yet expressive [20]. Other synthesis methods studied in isolation include additive and subtractive synthesis [9, 13, 7] and physical modeling [21, 22]. By isolation, we mean settings where the effect of individual parameters on the output sound remains tractable. For example, we exclude studies using commercial Virtual Studio Technology (VST) which can obscure the interactions between modules, losses, and outputs.
Finally, recent years have seen a growing interest in differentiable DSP (DDSP) [9], which integrates gradient-based optimization [23, 24] with DSP building blocks. Implementing complex DSP functions in a differentiable manner remains challenging, and robust differentiable audio similarity measures are still under active investigation [12, 14, 15].
II-C Sound Representation and Loss Functions
@@
$L(\theta,x_{0})=d(\phi(x_{\theta}),\phi(x_{0}))$
A proxy representation is the output of the function $\phi$, which can be thought of as a feature extraction function that maps the sounds $x_{0}$ and $x_{\theta}$ to their respective representations. The proportionality or distance metric $d$ has typically been the L1 or L2 distance [5, 28], with L1 being calculated as the mean of the absolute difference between every point in the proxy representation:
$L(\theta,x_{0})=\left\|\phi(x_{0})-\phi(x_{\theta})\right\|_{1}$
Here we discuss four methods of audio representation and the corresponding loss functions, including technical justifications for our novel methods.
II-C1 Parameter Loss
@@ II-C2 Fourier Spectrograms
Fourier-based transformations such as short-time Fourier transforms (STFT), Mel-spectrograms, and Mel-frequency cepstral coefficients have been viewed as the de facto and state-of-the-art representation of audio [29, 10, 30], however, there are many issues associated with their use in sound-matching [5, 14, 16, 15]. Fourier transformations allow for the conversion of a signal from the time-domain to the frequency domain. Audio spectrograms can be generated by segmentation of a piece of audio into overlapping windows followed by the application of Fourier transforms to each window. They are costly to compute, but provide a better temporal view of changes in frequency content [31, 26]. There are different types of spectrograms that have a basis in Fourier transformations, but the most notable and commonly used is the STFT. What we call Fourier-based Spectrograms are variations on the STFT approach. For example, Mel-Spectrograms bin frequencies on a near-logarithmic scale to better match human perception of frequencies [31]. Multi-scale spectrograms (MSS) used in recent works are a simple weighted average of multiple spectrograms with different parameters such as window size, number of frequency bins, and hop size [9, 14]; this may provide some improvements at a higher computational cost [5, 9].
A common limitation of spectrogram-based losses is their sensitivity to global gain: two signals with identical spectral shape but different overall amplitude can yield large errors under L1/L2 norms. Here we propose and test Scale-Invariant Mean Squared Error (SIMSE) as an alternative to L1/L2 norms in spectrogram comparisons. SIMSE normalizes the amplitudes before comparison and emphasizes proportional differences in spectral envelopes [32]. Perceptually, listeners are often tolerant of loudness changes while remaining sensitive to timbral shape, therefore this property could be advantageous for subtractive synthesis, where filter cutoffs reshape spectral balance without predictable changes in total energy. Although SIMSE has been applied in other contexts such as audio reconstruction, to our knowledge its use as a differentiable spectrogram loss in iterative sound-matching is novel.
II-C3 Joint-Time Frequency Spectrum
@@
Recent works have focused on the limitations of parameter and spectral loss functions in sound-matching [14, 15], seeking to create more effective general solutions for the comparison of audio. Noting the weaknesses of comparing STFT spectrograms (such as alignment and loudness distances), Vahidi et al. proposed differentiable Joint-Time Frequency Scattering (JTFS) [33] as an alternative to spectrogram loss in sound-matching, and showed improved performance in sound-matching with differentiable chirplet synthesizers [14]. JTFS is the result of the application of a 2D wavelet transformation to the time-frequency representation of a signal [33] and has been reported as more sensitive to mesostructural features such as melody, syncopation, and textural contrast [14].
II-C4 Dynamic Envelope Warping
@@
Dynamic Time Warping (DTW) is a method for measuring similarity between multi-dimensional time-series [34, 31, 35]. Given any two time-series $X=\{x_{1},x_{2},...,x_{m}\}$ and $Y=\{y_{1},y_{2},...,y_{n}\}$, we have indices $i\in\{1...m\}$ and $j\in\{1...n\}$ defining $X$ and $Y$. When the series are warped, these indices change to expand or contract different portions of the series. To borrow the notation given by Muller [31], warped indices are a sequence $p=(p_{1},...,p_{L})$, where $p_{\ell}=(m_{\ell},n_{\ell})\in[1:m]\times[1:n]\text{ for }\ell\in[1:L]$, meaning that the indices for $X$ and $Y$ are reorganized under special conditions. In classical DTW, these conditions are monotonicity, boundary matching, and single step-size. DTW measures the distance between the time-series after alignments, typically using Euclidean distance, such that the distance between a time-series and shifted versions of itself would be 0, regardless of shift amount [36]. Additional rules can be imposed to keep alignments locally constrained [37, 38].
DTW provides robustness to local temporal shifts, making it well suited for comparing modulated signals where the perceptual similarity lies in envelope dynamics rather than precise onset alignment. For example, two tremolo signals with slightly different phase are perceptually similar but would appear distant under spectrogram L1. A possible use case to consider is the application of DTW to amplitude envelopes, which may be able to match two sounds with similar rates of loudness modulation regardless of alignments. To our knowledge, DTW applied in this way has not been used as a loss in sound-matching (whether differentiable and iterative or not).
II-D In-Domain Versus Out-of-Domain
@@
The choice of domain depends on whether we want to use the same synthesizer for the target and output sounds, the scenario that is called in-domain, or have target sounds that came from sources other than the synthesizer, or out-of-domain. To paraphrase the description given by Masuda et al. [12], if $g$, the synthesizer of choice, can accurately replicate the target sound $x_{0}$, or put differently, if $x_{0}$ itself is an output of $g$, then the sound-matching task is in-domain. If $x_{0}$ is not an output of $g$, then the sound-matching task is out-of-domain. In-domain tasks in general are simpler, and often there is a guarantee that there is a correct answer to the sound-matching problem, particularly if the goal is accurate replication of the sound. If the target sound is out-of-domain, replication is not guaranteed, and the goal becomes the imitation of some aspect of sound.
@@
II-E Supervised Versus Direct Optimization
The choice of heuristics is yet another important attribute sound-matching. The goal of sound-matching is to find the optimal parameters $\theta^{*}$ that minimize the loss between the synthesizer output and the target sound.
$\theta^{*}=\arg\min_{\theta}L(\theta,x_{0})$
The heuristics (i.e., how $\theta^{*}$ is approximated) used in past works can be broadly split into the two categories of direct optimization and supervised (or inference) methods. Direct optimization refers to the iterative generation of a sound output, measurement of similarity between target and output, and application of updates to the parameters to maximize similarity (or minimize loss) [6, 10, 30, 14]; while supervised methods use large datasets of synthesizer sounds and corresponding parameters to learn the generation objective, commonly with the use of DNNs [9, 7, 30, 8]. These models often make their parameter predictions in a single step (or 1-shot).
Genetic algorithms (GA) [39] have been the earliest and most common heuristic for direct sound-matching [6, 10, 30]. These algorithms start with arbitrary parameter sets that can be treated as an evolving population where the genomes are the parameter values. The most fit members of the group are the parameters which perform the best in the loss function, and create a new generation of parameters via mutation (random change in subset of parameters) and crossovers (combination of parameter sets); this process repeats until the goal or a maximum number of generations is reached. Rather than using random mutations, differentiable approaches allow goal-oriented updates to the synthesizer parameters (the goal being the minimization of loss), but with some drawbacks: other than requiring more computation power, differentiable functions require careful implementation of operations that are continuous and numerically stable; this makes the implementation of signal processing functions quite difficult, contributing to the scarcity of works in this domain.